OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation 2x + 4y = 12.

OpenStudy (anonymous):

solve for y 4y= 12- 2x y = 3 -1/2x gradient = -1/2 intercept = the y value when x = 0 that is 3

OpenStudy (anonymous):

brig it in form of y=mx+c were m=slope c=y intercept 4y=-2x+12 y=-x/2 +3

OpenStudy (anonymous):

slope = -a/b and y intrcept is c so.. ans is -1/2 and 3